# spacing=nonfrench

### v1

@@ 2.1 Boundary Values of Functions $\mathcal{B}_{p,p}^{1/p}(\Omega)$
When dealing with a function $u$ in $\mathcal{B}_{p,p}^{1/p}(\Omega_{i})$ that is continuous up to the boundary the boundary values are obvious: it is just the restriction to the boundary, and moreover we have just seen that if we take the harmonic extension of this boundary value in $\Omega_{e}$ we get a function $v\in\mathcal{B}_{p,p}^{1/p}(\Omega_{e})$ with a norm equivalent to the one we started with. Now let us consider a general function $u\in\mathcal{B}_{p,p}^{1/p}(\Omega_{i})$: as we have seen, this function is the limit of a sequence $(u_{n})\in\mathcal{C}_{p}(\Omega_{i})$ and we put $f_{n}$ to be the boundary value of $u_{n}$. The harmonic extension to $\Omega_{e}$ of $f_{n}$ is then a Cauchy sequence in $\mathcal{B}_{p,p}^{1/p}(\Omega_{e})$, thus convergent to some $v\in\mathcal{B}_{p,p}^{1/p}(\Omega_{e})$ and we may regard the couple $(u,v)$ as an abstract version of the boundary value of $u$, and also $v$.
@@
If we want to be more concrete, we use the Riemann maps $\varphi_{i,e}$: $u\circ\varphi_{i}\in\mathcal{B}_{p,p}^{1/p}(\mathbb{D}_{i})\subset h^{p}(\mathbb{D})$, the classical harmonic Hardy space of the disk, so it has radial boundary values almost everywhere on the circle defining a function $b_{i}\in L^{p}(\mathbb{T})$. Notice that this fact translates in $\Omega_{i}$ by saying that $u$ has limits along $\omega_{i}$-almost every internal ray, where $\omega_{i}$ stands for harmonic measure in $\Omega_{i}$. Notice that this limit $b_{i}$ characterizes $u$. We may do the same thing with $v$ and define an analogous function $b_{e}\in L^{p}(\mathbb{T})$. When $u\in\mathcal{C}_{p}(\Omega_{i})$ we have seen that $b_{e}=V_{h}(b_{i})$ where $h$ is the conformal welding of $\Gamma$. Unfortunately it does not make sense in $L^{p}(\mathbb{T})$ since $h$ need not be absolutely continuous with respect to the Lebesgue measure, a fact which transfers geometrically to the fact that $\omega_{i}$ and $\omega_{e}$ need not be mutually absolutely continuous. Nevertheless, the relation $b_{e}=V_{h}(b_{i})$ remains true if we interpret $V_{h}$ on $\mathcal{B}_{p,p}^{1/p}(\mathbb{T})$ as the completion of $V_{h}|_{\mathcal{C}_{p}^{1/p}(\mathbb{T})}$.
@@
In the case of $p=2$, Schippers-Staubach [36] has given an even more concrete description of the boundary values: we outline their argument. On the one hand, for any $u\in\mathcal{B}_{2,2}^{1/2}(\Omega_{i})$ we have $u\circ\varphi_{i}\in\mathcal{B}_{2,2}^{1/2}(\mathbb{D}_{i})$ as before. In this case, $u\circ\varphi_{i}$ has radial boundary values everywhere on the circle $\mathbb{T}$ except on a Borel set $F_{1}$ of Logarithmic capacity zero, defining a function $b_{i}\in L^{2}(\mathbb{T})$. On the other hand, let $h:=\varphi_{i}^{-1}\circ\varphi_{e}$, the conformal welding with respect to the quasicircle $\Gamma$, that is a quasisymmetry (so is $h^{-1}$) as we pointed out above. Since a quasisymmetry takes a Borel set of Logarithmic capacity zero to a Borel set of Logarithmic capacity zero (see [2]) we see that $F_{2}:=h^{-1}(F_{1})$ is also a Borel set of Logarithmic capacity zero. Then $b_{e}=b_{i}\circ\varphi_{i}^{-1}\circ\varphi_{e}=V_{h}(b_{i})$ is well-defined on $\mathbb{T}\setminus F_{2}$ and belongs to $L^{2}(\mathbb{T})$. It is known that the union $F_{1}\cup F_{2}$ is also of Logarithmic capacity zero, and in particular, Lebesgue measure zero since a set of Logarithmic capacity zero has zero Lebesgue measure. Put $P_{e}(b_{e})$ to be the harmonic extension in $\mathbb{D}_{e}$ of $b_{e}$, and $v:=P_{e}(b_{e})\circ\varphi_{e}^{-1}$ is the harmonic extension in $\Omega_{e}$ of $b_{i}\circ\varphi_{i}^{-1}$. We can see that $u$ and $v$ have the same boundary values along every internal ray on $\Gamma$, except on a set whose harmonic measure is zero with respect to both $\Omega_{i}$ and $\Omega_{e}$. Moreover, since the spaces $\mathcal{B}_{p,p}^{1/p}(\Omega_{i,e})$ increase with $p$ the same argument also holds for the case of $1<p<2$.
@@ 2.2 Plemelj-Calderón Problem for $\mathcal{B}_{2,2}^{1/2}(\Gamma)$
We can now address Plemelj-Calderón problem for $\mathcal{B}_{2,2}^{1/2}(\Gamma)$. The argument is broken into two steps. In the first we consider the function in the space $D(\Gamma)$, the restriction space to $\Gamma$ of $C_{c}^{\infty}(\mathbb{C})$. The second step consists in an approximation process.
For any test function $f$, it holds that
@@
It is easily seen that $\widehat{F_{i}}$ and $\widehat{F_{e}}$ are holomorphic functions in $\Omega_{i}$ and $\Omega_{e}$, respectively with $\widehat{F_{e}}(z)=O(\frac{1}{|z|})$ at $\infty$. Furthermore, it can be shown that both $\widehat{F_{i}}$ and $\widehat{F_{e}}$ define continuous functions at each $z\in\mathbb{C}$. It follows from (2.4) that on $\Gamma$,
@@
Here, we have used that the quasicircle $\Gamma$ has zero area with respect to the two-dimensional Lebesgue measure. Taking derivative we have that
@@
Here, $\chi_{\Omega_{e}}$ is the characteristic function of the domain $\Omega_{e}$ and $B$ denotes the Beurling operator, i.e., the convolution with $\frac{1}{\pi}\text{p.v.}\frac{1}{z^{2}}$ which is an isometry of $L^{2}(\mathbb{C})$. Then,
@@
$\leq\|B(\bar{\partial}f\chi_{\Omega_{e}})\|_{L^{2}(\mathbb{C})}\leq\|\bar{\partial}f\|_{L^{2}(\Omega_{e})}\leq\|\bar{\partial}f\|_{L^{2}(\mathbb{C})}.$
Similarly, it follows from (2.6) that
@@
For any $f\in C_{c}^{\infty}(\mathbb{C})$, the integrals in (2.5) and (2.6) depend only on the boundary values $f$ on $\Gamma$, not on the specific extensions belonging to $B_{2,2}^{1/2}(\Omega_{i,e})$ of $f$ to respectively $\Omega_{i}$ and $\Omega_{e}$. In particular, $\widehat{F_{i}}$ and $\widehat{F_{e}}$ defined in respectively (2.5) and (2.6) equal to
@@
on $\Gamma$. Here, $u_{i}$ and $u_{e}$ are harmonic extensions of $f$ to respectively $\Omega_{i}$ and $\Omega_{e}$.
@@
Note that $u_{i}$ and $u_{e}$ are continuous on $\overline{\Omega}_{i}$ and $\overline{\Omega}_{e}$, respectively since the boundary function $f$ is continuous on $\Gamma$. By the Dirichlet principle, $u_{i,e}\in\mathcal{B}_{2,2}^{1/2}(\Omega_{i,e})$. Using the similar computation we can see that the holomorphic functions $F_{i}\in\mathcal{B}_{2,2}^{1/2}(\Omega_{i})$ and $F_{e}\in\mathcal{B}_{2,2}^{1/2}(\Omega_{e})$.
By the theorem of Gol’dshtein-Latfullin-Vodop’yanov [18] (see also [22]) $u_{i}$ can be extended to $U_{i}\in W^{1,2}(\mathbb{C})$, defining $U_{i}=u_{i}\circ R$ on $\Omega_{e}$ where $R$ is a quasiconformal reflection across $\Gamma$. There exists a sequence $(U_{i,n})$ of $C_{c}^{\infty}(\mathbb{C})$ converging to $U_{i}$ in $W^{1,2}(\mathbb{C})$ (see e.g. Ch.11 in [26]). Denote $U_{i,n}|_{\Gamma}$ by $f_{i,n}$. By what proceeds, $f_{i,n}$ may be written as $\widehat{F_{i,n}}-\widehat{F_{e,n}}$ on $\Gamma$. Here,
@@
Notice that $\widehat{F_{i,n}}$ and $\widehat{F_{e,n}}$ are holomorphic in $\Omega_{i},\,\Omega_{e}$ respectively, continuous in $\mathbb{C}$, and $F_{e,n}(z)=O(\frac{1}{|z|})$ as $z\to\infty$. Similar to the reasoning in (2.8) we have that
@@
where $P_{i}(f_{i,n})$ denotes the harmonic extension in $\Omega_{i}$ of $f_{i,n}$. We thus have that $f_{i,n}(z)\to f(z)$ at each point $z\in\Gamma$.
@@
By uniqueness of the decomposition, proved in [27] mainly based on an identity in [39], we see that these two decomposition (2.13),(2.14) are same as (2.7) so that $\widehat{F_{i}}=F_{i}$,  $\widehat{F_{e}}=F_{e}$ and $f=F_{i}-F_{e}$ on $\Gamma$. ∎
Concerning the above proof, we give two side remarks. Firstly, using a part of the argument in the proof of Lemma 2.2, we can see that $C_{c}^{\infty}(\mathbb{C})|_{\Gamma}$ is a dense subset of $\mathcal{B}_{2,2}^{1/2}(\Gamma)$. Precisely, for any $f\in\mathcal{B}_{2,2}^{1/2}(\Gamma)$, its harmonic extension $u_{i}$ in $\Omega_{i}$ can be extended to $U_{i}\in W^{1,2}(\mathbb{C})$. There exists a sequence $(U_{i,n})$ of $C_{c}^{\infty}(\mathbb{C})$ converging to $U_{i}$ in $W^{1,2}(\mathbb{C})$. Using the Dirichlet principle, we may conclude the assertion.
Secondly, we explain that if the Jordan curve $\Gamma$ is rectifiable then the assertion in Lemma 2.2 can be derived easily from the generalized Green’s formula (see p.150 in [28]). It says that let $f$ have $L^{1}$-derivatives in the domain $G$ and if $\overline{\Omega}\subset G$ is a Jordan domain with rectifiable boundary $\Gamma=\partial\Omega$ then
@@
Let $f\in C_{c}^{\infty}(\mathbb{C})$ and $\Gamma$ be a rectifiable Jordan curve and let $u_{e}$ be the harmonic extension of $f$, restricted on $\Gamma$, to $\Omega_{e}$, that is continuous on $\Omega_{e}\cup\Gamma$. By the Dirichlet principle, $u_{e}\in W^{1,2}(\Omega_{e})$. As we noticed before, $u_{e}$ can be extended to $U_{e}\in W^{1,2}(\mathbb{C})$ using the quasiconformal reflection across $\Gamma$. Note that $f-U_{e}$ has $L^{2}$-derivatives in $\mathbb{C}$. By the generalized Green’s formula,
@@
which equals to 0 since $(f-u_{e})|_{\Gamma}=0$ on $\Gamma$. Thus, we see that
@@
Here and in what follows, the notation $A\simeq B$" means there exists a constant $C$ such that $A/C\leq B\leq CA$. In the above the implicit constant depends only on $\Gamma$.

### v2

@@ 2.2 Boundary values of functions in $\mathcal{B}_{p,p}^{1/p}(\Omega)$
When dealing with a function $u$ in $\mathcal{B}_{p,p}^{1/p}(\Omega_{i})$ that is continuous up to the boundary, the boundary values are obvious: it is just the restriction to the boundary, and moreover we have just seen that if we take the harmonic extension of this boundary value in $\Omega_{e}$ we get a function $v\in\mathcal{B}_{p,p}^{1/p}(\Omega_{e})$ with a norm equivalent to the one we started with. Now let us consider a general function $u\in\mathcal{B}_{p,p}^{1/p}(\Omega_{i})$: as we have seen, this function is the limit of a sequence $(u_{n})\in\mathcal{C}_{p}(\Omega_{i})$ and we put $f_{n}$ to be the boundary value of $u_{n}$. The harmonic extension to $\Omega_{e}$ of $f_{n}$ is then a Cauchy sequence in $\mathcal{B}_{p,p}^{1/p}(\Omega_{e})$, thus convergent to some $v\in\mathcal{B}_{p,p}^{1/p}(\Omega_{e})$ and we may regard the couple $(u,v)$ as an abstract version of the boundary value of $u$, and also $v$. The set of such boundary functions is denoted by $\mathcal{B}_{p,p}^{1/p}(\Gamma)$.
@@
If we want to be concrete, we use the Riemann maps $\varphi_{i,e}$: $u\circ\varphi_{i}\in\mathcal{B}_{p,p}^{1/p}(\mathbb{D}_{i})\subset h^{p}(\mathbb{D})$, the classical harmonic Hardy space of the disk, so it has radial boundary values almost everywhere on the circle defining a function $b_{i}\in L^{p}(\mathbb{T})$. Notice that this fact translates in $\Omega_{i}$ by saying that $u$ has limits along $\omega_{i}$-almost every internal ray, where $\omega_{i}$ stands for harmonic measure in $\Omega_{i}$. Notice that this limit $b_{i}$ characterizes $u$. We may do the same thing with $v$ and define an analogous function $b_{e}\in L^{p}(\mathbb{T})$. When $u\in\mathcal{C}_{p}(\Omega_{i})$ we have seen that $b_{e}=V_{h}(b_{i})$ where $h$ is the conformal welding of $\Gamma$. Unfortunately, it does not make sense in $L^{p}(\mathbb{T})$ since $h$ need not be absolutely continuous with respect to the Lebesgue measure, a fact that geometrically transfers to the fact that $\omega_{i}$ and $\omega_{e}$ need not be mutually absolutely continuous. Nevertheless, the relation $b_{e}=V_{h}(b_{i})$ remains true if we interpret $V_{h}$ on $\mathcal{B}_{p,p}^{1/p}(\mathbb{T})$ as the completion of $V_{h}|_{\mathcal{C}_{p}^{1/p}(\mathbb{T})}$.
@@
In the case of $p=2$, Schippers-Staubach ([40], [43]) has given an even more concrete description of the boundary values: we outline their argument. On the one hand, for any $u\in\mathcal{B}_{2,2}^{1/2}(\Omega_{i})$ we have $u\circ\varphi_{i}\in\mathcal{B}_{2,2}^{1/2}(\mathbb{D}_{i})$ as before. In this case, $u\circ\varphi_{i}$ has radial boundary values everywhere on the circle $\mathbb{T}$ except for a Borel set $F_{1}$ of Logarithmic capacity zero, defining a function $b_{i}\in L^{2}(\mathbb{T})$. On the other hand, let $h=\varphi_{i}^{-1}\circ\varphi_{e}$, the conformal welding with respect to the quasicircle $\Gamma$, that is a quasisymmetry (so is $h^{-1}$) as we pointed out above. Since a quasisymmetry takes a Borel set of Logarithmic capacity zero to a Borel set of Logarithmic capacity zero (see [2]) we see that $F_{2}:=h^{-1}(F_{1})$ is also a Borel set of Logarithmic capacity zero. Then $b_{e}=b_{i}\circ\varphi_{i}^{-1}\circ\varphi_{e}=V_{h}(b_{i})$ is well-defined on $\mathbb{T}\setminus F_{2}$ and belongs to $L^{2}(\mathbb{T})$. It is known that the union $F_{1}\cup F_{2}$ is also of Logarithmic capacity zero, and, in particular, Lebesgue measure zero since a set of Logarithmic capacity zero has zero Lebesgue measure. Put $P_{e}(b_{e})$ as the harmonic extension in $\mathbb{D}_{e}$ of $b_{e}$, and $v:=P_{e}(b_{e})\circ\varphi_{e}^{-1}$ is the harmonic extension in $\Omega_{e}$ of $b_{i}\circ\varphi_{i}^{-1}$. We can see that $u$ and $v$ have the same boundary values along every internal ray on $\Gamma$, except on a set whose harmonic measure is zero with respect to both $\Omega_{i}$ and $\Omega_{e}$. Moreover, since spaces $\mathcal{B}_{p,p}^{1/p}(\Omega_{i,e})$ increase with $p$, the same argument holds for the case of $1<p<2$.
@@ 2.3 Plemelj-Calderón problem for $\mathcal{B}_{2,2}^{1/2}(\Gamma)$
We can now address Plemelj-Calderón problem for $\mathcal{B}_{2,2}^{1/2}(\Gamma)$. The main result of this subsection is Theorem 2.3, which has been proven in the literature; see [40], [29]. However, the proof we present below is also worth recording, and will be used further in later subsections. The argument is broken down into two steps. In the first step, we consider the function in the space $C_{c}^{\infty}(\mathbb{C})|_{\Gamma}$, the restriction space to $\Gamma$ of $C_{c}^{\infty}(\mathbb{C})$. The second step consists in an approximation process.
For any test function $f$, we have the following.
@@
It is easily seen that $\widehat{F_{i}}$ and $\widehat{F_{e}}$ are holomorphic functions in $\Omega_{i}$ and $\Omega_{e}$ respectively with $\widehat{F_{e}}(z)=O(\frac{1}{|z|})$ at $\infty$. Furthermore, it can be shown that both $\widehat{F_{i}}$ and $\widehat{F_{e}}$ define continuous functions in $\mathbb{C}$. It follows from (2.4) that on $\Gamma$,
@@
Here, we have used the fact that the quasicircle $\Gamma$ has zero area with respect to the two-dimensional Lebesgue measure. Taking the derivative on both sides of (2.5) we get
@@
Here, $\chi_{\Omega_{e}}$ is the characteristic function of the domain $\Omega_{e}$ and $B$ denotes the Beurling operator, i.e., the convolution with $\frac{1}{\pi}\text{p.v.}\frac{1}{z^{2}}$ which is an isometry of $L^{2}(\mathbb{C})$. Then
@@
$\leq\|B(\bar{\partial}f\chi_{\Omega_{e}})\|_{L^{2}(\mathbb{C})}=\|\bar{\partial}f\|_{L^{2}(\Omega_{e})}\leq\|\bar{\partial}f\|_{L^{2}(\mathbb{C})}.$
Similarly, from (2.6) it follows that
@@
For any $f\in C_{c}^{\infty}(\mathbb{C})$, the integrals in (2.5) and (2.6) depend only on the boundary values $f$ on the quasicircle $\Gamma$, not on the specific extensions belonging to $B_{2,2}^{1/2}(\Omega_{i,e})$ of $f$ to $\Omega_{i}$ and $\Omega_{e}$. In particular, $\widehat{F_{i}}$ and $\widehat{F_{e}}$ defined in respectively (2.5) and (2.6) equal to
@@
on $\Gamma$. Here, $u_{i}$ and $u_{e}$ are harmonic extensions of $f$ to $\Omega_{i}$ and $\Omega_{e}$.
@@
Note that $u_{i}$ and $u_{e}$ are continuous on $\overline{\Omega}_{i}$ and $\overline{\Omega}_{e}$, respectively, since the boundary function $f$ is continuous on $\Gamma$. By the Dirichlet principle, $u_{i,e}\in\mathcal{B}_{2,2}^{1/2}(\Omega_{i,e})$. Using a similar computation to that of $\widehat{F_{i}}$ and $\widehat{F_{e}}$, we can see that the holomorphic functions $F_{i}\in\mathcal{B}_{2,2}^{1/2}(\Omega_{i})$ and $F_{e}\in\mathcal{B}_{2,2}^{1/2}(\Omega_{e})$.
By the theorem of Gol’dshtein-Latfullin-Vodop’yanov [18] (see also [22]) saying that a simply connected domain $\Omega$ satisfies the condition $W^{1,2}(\mathbb{C})|_{\Omega}=W^{1,2}(\Omega)$ if and only if $\partial\Omega$ is a quasicircle, $u_{i}$ can be extended to $U_{i}\in W^{1,2}(\mathbb{C})$ by defining $U_{i}=u_{i}\circ R$ on $\Omega_{e}$ where $R$ is a quasiconformal reflection across $\Gamma$. There exists a sequence $(U_{i,n})$ of $C_{c}^{\infty}(\mathbb{C})$ converging to $U_{i}$ in $W^{1,2}(\mathbb{C})$ (see, e.g., Ch.11 in [26]). Denote $U_{i,n}|_{\Gamma}$ by $f_{i,n}$. By what proceeds, $f_{i,n}$ may be written as $\widehat{F_{i,n}}-\widehat{F_{e,n}}$ on $\Gamma$. Here,
@@
Notice that $\widehat{F_{i,n}}$ and $\widehat{F_{e,n}}$ are holomorphic in $\Omega_{i},\,\Omega_{e}$, respectively, continuous in $\mathbb{C}$, and $\widehat{F_{e,n}}(z)=O(\frac{1}{|z|})$ as $z\to\infty$. Similarly to the reasoning in (2.8) we get
@@
where $P_{i}(f_{i,n})$ denotes the harmonic extension in $\Omega_{i}$ of $f_{i,n}$. Thus, we have $f_{i,n}(z)\to f(z)$ at each point $z\in\Gamma$.
@@
By the uniqueness of the decomposition (see e.g. [28]), we see that these two decompositions (2.13),(2.14) are the same as (2.7) so that $\widehat{F_{i}}=F_{i}$,  $\widehat{F_{e}}=F_{e}$ and $f=F_{i}-F_{e}$ on $\Gamma$. ∎
Concerning the above proof, we make two side remarks. Firstly, using a part of the argument in the proof of Lemma 2.2, we can see that $C_{c}^{\infty}(\mathbb{C})|_{\Gamma}$ is a dense subset of $\mathcal{B}_{2,2}^{1/2}(\Gamma)$. Specifically, for any $f\in\mathcal{B}_{2,2}^{1/2}(\Gamma)$, its harmonic extension $u_{i}$ in $\Omega_{i}$ can be extended to $U_{i}\in W^{1,2}(\mathbb{C})$. There exists a sequence $(U_{i,n})$ of $C_{c}^{\infty}(\mathbb{C})$ converging to $U_{i}$ in $W^{1,2}(\mathbb{C})$. Using the Dirichlet principle, we may conclude the assertion.
Secondly, we explain that if the quasicircle $\Gamma$ is rectifiable then the assertion in Lemma 2.2 can be derived easily from the generalized Green’s formula (see p.150 in [30]). It says that if $f$ has $L^{1}$-derivatives in the domain $G$ and if $\overline{\Omega}\subset G$ is a Jordan domain with rectifiable boundary $\Gamma=\partial\Omega$ then
@@
Let $f\in C_{c}^{\infty}(\mathbb{C})$ and $\Gamma$ be a rectifiable Jordan curve and let $u_{e}$ be the harmonic extension of $f$, restricted on $\Gamma$, to $\Omega_{e}$, which is continuous in $\Omega_{e}\cup\Gamma$. By the Dirichlet principle, $u_{e}\in W^{1,2}(\Omega_{e})$. As we have noticed before, $u_{e}$ can be extended to $U_{e}\in W^{1,2}(\mathbb{C})$ using the quasiconformal reflection across $\Gamma$. Note that $f-U_{e}$ has $L^{2}$-derivatives in $\mathbb{C}$. By the generalized Green’s formula,
@@
which equals 0 since $(f-u_{e})|_{\Gamma}=0$ on $\Gamma$. Thus, we see that
@@
Here and in what follows, the notation $A\simeq B$" means that there exists a constant $C$ such that $A/C\leq B\leq CA$. In the above, the implicit constant depends only on $\Gamma$.